OpenStudy (anonymous):

Identify the graph of 3y – 5x ≥ –6

OpenStudy (anonymous):

3y - 5x >= -6 3y >= 5x - 6 y >= 5/3x - 2 y intercept is -2 x intercept is found by subbing in 0 for y 3(0) - 5x >= -6 -5x >= -6 x >= 6/5 or 1 1/5 so plot a point on (0,-2) and (6/5,0) -- if it helps, 6/5 = 1 1/5 connect your points with a solid line because of the equal sign, and shade above the line because it is greater then.